Project

General

Profile

NDOS Commissioning Meeting 2011/07/14

Attendance: Jon, Peter, Chad, Susan, Gary, Raphael, Mark Mathis, Mat, Denis, Mark, Andrew, Minerba, Holger, Leon, Sue, Jan, Nick, Dominic, Zeb, Gavin, Martin,

Agenda:

  1. NDOS Update
  2. DataCheck Update
  3. DCS Update
  4. DAQ Update
  5. AOB

NDOS Update

*
  • APD recovery (Leon): working on modifying test stand to efficiently bias APDs for long periods of time. Hope to ship some APDs back next week.
  • Sue: Ken is concerned about stressing seals on APDs by frequent enabling and disabling cooling.
  • Mat placed balloons around optical connectors on 5 modules on diblock 3 and 5 failed golden channels.
  • Leon: pressure tests should be done on modules
  • Mark: sucking on optical connection may be better way to test for tightness.
  • Consecutive DSO scans were done yesterday
  • 2 minute turn off, we lost 1/75
  • 10 minute turn off, we lost 20%

DCS Update

  • Martin:
    • Occassionally CSS was not able to see all DCMs
    • Increased amount of memory that CSS can use; it is now working much more smoothly
    • Overview page is now part of base release (default)
    • Jon: documentation? Martin: wrote some basic info.
    • Gennadiy is working on alarmer and making good process
  • Chad: hasn't seen any real improvement in NDSB since 2nd humidifier was installed. Hovering around 40% RH. This weekend it should become much more humid, so we'll keep an eye on it.
  • Gary: what's the point of recording env. data in online form if shifter doesn't know what to look for?
  • Jon: we'll work on this

DAQ Update

  • Peter: nothing relevant to report

DataCheck Update

AOB